# Spectrum of high-frequency gravitational waves from graviton bremsstrahlung by the decay of inflaton: case with polynomial potential

### v1

@@ 2.1 The model
Here $M_{p}=1/\sqrt{8\pi G_{N}}\simeq 2.4\times 10^{18}$eV the reduced Planck mass.
@@ 3.1 Bosonic reheating
where $\frac{d\Gamma^{1\to 3}_{\phi_{n}}}{dE_{\omega}}$ is the three-body decay rate of a massive particle having a mass $n\omega$ computed based on the standard QFT. Remarkably, the same weight $b_{n}$ as in Eq. (3.3) appears in the above decay rate. Thus, the graviton bremsstrahlung can be regarded as a process of decays of massive particles for which particles labeled with an integer $n$ have mass $n\omega$, constitute a fraction $b_{n}$, and decay with the rate $\frac{d\Gamma^{1\to 3}_{\phi_{n}}}{dE_{\omega}}$. Nevertheless, in order to make a comparison between our results and those in the literature, in what follows, we will retain the original meaning of the particle picture.
@@ 3.2 Fermionic reheating
where $\frac{d\Gamma^{1\to 3}_{n}}{dE_{\omega}}$ is the three-body decay rate of a particle with mass $n\omega$ computed by the standard QFT;
@@ 5.2 Fermionic reheating
Figure 6: GW spectrum based on the particle picture (left panel) and the classical field picture (right panel) in the case where inflaton decays to fermions. The exponent $k$ is from $k=2$ to $k=10$, $\lambda=10^{-11}$ and $y=0.5$. Solid orange lines represent the sensitivities of future gravitational wave detectors.
@@ Acknowledgments
This work was supported by the JSPS KAKENHI Grant Number JP23K03411 (TS).
@@ C.1 Bosonic reheating
$\left.\frac{d\rho_{\text{GW}}}{dE_{\omega}}\right|_{a=a_{\text{rh}}}\simeq\rho_{\phi}(a_{\text{rh}})\frac{1}{\Gamma_{\phi}^{1\to 2}}\left.\frac{d\Gamma^{1\to 3}}{dE_{\omega}}\right|_{a=a_{\text{rh}}}.$ (C.15)

### v2

@@ 2.1 The model
Here $M_{p}=1/\sqrt{8\pi G_{N}}\simeq 2.4\times 10^{18}$ GeV is the reduced Planck mass, and $\lambda$ is a constant that parametrizes the energy scale of the potential after inflation. Since the potential associated with the inflationary phase may differ from the post-inflationary potential, the value of $\lambda$ can only be determined from CMB observations, such as those by the PLANCK mission, once the inflationary potential is specified. However, because the specific form of the inflationary potential is not directly relevant to our analysis (i.e., the assumption in Eq. (2.2) suffices), we refrain from adopting a specific potential. Instead, we simply use $\lambda=10^{-11}$ as a representative value in our numerical calculations. In Sec. 5, we provide an approximate relation that describes how the GW spectrum changes with variations in $\lambda$. This relation enables a straightforward determination of the GW spectrum for different values of $\lambda$.
@@ 3.1 Bosonic reheating
where $\frac{d\Gamma^{1\to 3}_{\phi_{n}}}{dE_{\omega}}$ is the three-body decay rate of a massive particle having a mass $n\omega$ computed based on the standard QFT,
$\frac{d\Gamma_{\phi}^{1\to 3}}{dE_{\omega}}=\frac{1}{32\,\pi^{3}}\left(\frac{\mu}{M_{P}}\right)^{2}\frac{(1-2x_{n})^{2}}{4x_{n}}\,,~{}~{}~{}~{}~{}\quad 0<x_{n}<\frac{1}{2}.$ (3.10)
Remarkably, the same weight $b_{n}$ as in Eq. (3.3) appears in the above decay rate. Thus, the graviton bremsstrahlung can be regarded as a process of decays of massive particles for which particles labeled with an integer $n$ have mass $n\omega$, constitute a fraction $b_{n}$, and decay with the rate $\frac{d\Gamma^{1\to 3}_{\phi_{n}}}{dE_{\omega}}$. Nevertheless, in order to make a comparison between our results and those in the literature, in what follows, we will retain the original meaning of the particle picture.
@@ 3.2 Fermionic reheating
where $\frac{d\Gamma^{1\to 3}_{\phi_{n}}}{dE_{\omega}}$ is the three-body decay rate of a particle with mass $n\omega$ computed by the standard QFT,
@@ 5.1 Bosonic reheating
In the same manner as we have derived the scaling of the GW spectrum with the coupling constant, the dependence of the GW spectrum on the parameter $\lambda$ can also be derived. First, it can be shown that the peak frequency of GW spectrum scales with the parameter $\lambda$ as
$f_{\mathrm{GW},0}\propto\begin{cases}\lambda^{\frac{1}{4k-4}}&(k\leq 6)\,,\\ \lambda^{\frac{4k-13}{12(k-1)}}&(k\geq 8)\,.\end{cases}$ (5.5)
The height of the GW spectrum at the peak frequency depends on the parameter $\lambda$ as
$\Omega_{\rm GW}^{0}\propto\begin{cases}\lambda^{\frac{1}{k-1}}&(k\leq 6)\,,\\ \lambda^{\frac{k-4}{3k-3}}&(k\geq 8)\,.\end{cases}$ (5.6)
Furthermore, in the case of $k=2$, the inflaton may undergo fragmentation through a process of tachyonic resonance when coupled to a boson via a trilinear interaction [44]. Specifically, the trilinear coupling $\mu\phi b^{2}$ contributes a tachyonic squared mass term of the form $m_{b}^{2}\sim\mu\phi$ to the boson. This squared mass becomes negative during certain phases of $\phi$’s oscillation, leading to tachyonic resonance. However, when considering the self-interaction of bosonic fields, represented by the term $\lambda_{b}b^{4}$, a positive squared mass term $\lambda_{b}\langle b^{2}\rangle$ is introduced. This backreaction suppresses the tachyonic effects [44, 6, 12]. In our study, we assume the absence of both the trilinear and self-interactions.
@@ 5.2 Fermionic reheating
Figure 6: GW spectrum based on the particle picture (left panel) and the classical field picture (right panel) in the case where inflaton decays to fermions. The exponent $k$ is from $k=2$ to $k=10$, $\lambda=10^{-11}$ and $y=0.01$. Solid orange lines represent the sensitivities of future gravitational wave detectors.
@@
The dependence of the GW spectrum on the potential parameter $\lambda$ can also be derived as bosonic case. First, the peak frequency of GW spectrum scales with the parameter $\lambda$
$f_{\mathrm{GW},0}\propto\lambda^{\frac{1}{4}}$ (5.11)
The height of the GW spectrum at the peak frequency depends on the parameter $\lambda$ as
$\Omega_{\rm GW}^{0}\propto\lambda$ (5.12)
@@ 5.3 Discussions
For cases where $k>2$, it is known that the inflaton condensate undergoes fragmentation due to self-interactions, resulting in inhomogeneities in the inflaton [46, 47, 48, 49]. Previous studies have shown that this fragmentation typically occurs more than 5 e-folds after the end of inflation. As discussed in Appendix C, for $k>2$, GWs from graviton bremsstrahlung are dominantly produced near the beginning of the oscillatory phase, except in the cases of bosonic reheating with $k=4$ and $k=6$. Therefore, if such bosonic decay is not the main channel, the final GW spectrum derived in this paper remains unchanged, even when accounting for the effects of fragmentation. For bosonic reheating with $k=4$ and $k=6$, however, GWs are dominantly generated around the time of reheating. Even in such scenarios, if inflaton fragmentation occurs after reheating, its impact on the GW spectrum will be not significant. According to [47], self-fragmentation is completed by the inflaton quanta produced within the first narrow instability band described by the Floquet theorem. This provides an estimate for the time at which backreaction from fluctuations becomes significant. Using the results of this paper and imposing the condition that the reheating time is less than the backreaction time, a lower bound on the coupling constant $\mu$ can be derived. For instance, assuming $\lambda=10^{-11}$, we find that the bound is given by $\mu>2.87\times 10^{10}\,\text{GeV}$ for $k=4$ and $\mu>1.29\times 10^{6}\,\text{GeV}$ for $k=6$, respectively. If $\mu$ falls below these thresholds, the effects of fragmentation on the final GW spectrum will become significant. Computing the GW spectra including the effects of fragmentation is beyond the scope of this paper.
Moreover, in this study, gravitational reheating is assumed to be negligible. Previous research on gravitational reheating has demonstrated that for $w_{\phi}\gtrsim 0.65$ [50, 51, 52, 53, 54], corresponding to $k>9$, gravitational reheating can independently reheat the universe and satisfy the temperature constraints imposed by BBN. As $k$ increases further, gravitational reheating becomes a non-negligible effect. Therefore, this study does not consider potentials with $k>10$.
@@ Acknowledgments
We would like to thank Drazen Glavan for giving us useful comments. This work was supported by the JSPS KAKENHI Grant Number JP23K03411 (TS). Y.J. was supported by JST SPRING, Japan Grant Number JPMJSP2180.
@@ C.1 Bosonic reheating
$\left.\frac{d\rho_{\text{GW}}}{dE_{\omega}}\right|_{a=a_{\text{rh}}}\simeq\rho_{\phi}(a_{\text{rh}})\frac{1}{\Gamma_{\phi}^{1\to 2}}\left.\frac{d\Gamma_{\phi}^{1\to 3}}{dE_{\omega}}\right|_{a=a_{\text{rh}}}.$ (C.15)